OpenStudy (anonymous):

Estimate the sum of 12.43, 17.47, 18.63 and 6.39 by each of the following methods: (a) Rounding off the numbers to the nearest integers (b) Using the compensation strategy.

OpenStudy (anonymous):

For (a), I did it in this way Sum = 12 + 17 + 19 + 6 = 54 What about (b), how should I do it?

OpenStudy (anonymous):

in a, what happened to the the 47?

OpenStudy (anonymous):

Sorry.. typo. 17.47

OpenStudy (anonymous):

12 + 17 + 19 + 6 = (12-2) + (17+3) + (19+1) + (6-2) ??? tell you the truth, i've never heard of the compensation strategy...:(

OpenStudy (anonymous):

Nope :( I don't like estimation!!!!

OpenStudy (anonymous):

The answer of (b) is different from (a), according to the book.

OpenStudy (anonymous):

sorry.... idk how to do it... i'm not familiar with the compensation thingy...

OpenStudy (anonymous):

It's okay. Thanks for trying!!!

OpenStudy (anonymous):

for b i think its something like this 10+ 20 + 20+ 5 -1 =54

OpenStudy (anonymous):

i think thats how it is, you round the numbers to the closest five or tens then, you add up what you subtracted/added from each to the end

OpenStudy (anonymous):

Nope. It's not 54...

OpenStudy (anonymous):

a) 12 + 18 +19 +6 b) 10 + 20 +20 +5 =55?

OpenStudy (anonymous):

For (a), I think it's 12 + 17 + 19 + 6 =54 For (b), you've got the answer, but why did you do in that way?

OpenStudy (anonymous):

b says use the compensation strategy

OpenStudy (anonymous):

Yes?!

OpenStudy (anonymous):

But that compensation strategy exactly is?

OpenStudy (anonymous):

*that -> what

OpenStudy (anonymous):

this is where i got it http://wiki.answers.com/Q/How_do_you_do_math_compensation_subtraction

OpenStudy (anonymous):

i'm not sure if that is exactly what your book asks for

OpenStudy (anonymous):

But you're getting the exact answer. In this case, it's not.

OpenStudy (anonymous):

Another interesting thing is that, why does it compensate the last one, but not the first one - i mean both of them together.

OpenStudy (anonymous):

here's another good explanation, scroll to the 6th page http://www.coolclass.ca/pdf/Estimation.pdf

OpenStudy (anonymous):

I'm not sure what you mean

OpenStudy (anonymous):

Hmm... like the questions there are all integers, not decimal number. When we round the number 12.43, 17.47, 18.63 and 6.39 12 17 19 6 Compensate: 10 (+2) 20 (-3) 20 (-1) 5(+1) => 10 + 20 + 20 + 5 => 55 .... like that?

OpenStudy (anonymous):

looks about right

OpenStudy (anonymous):

well i hoped that helped

OpenStudy (anonymous):

But why would 17 change to 20? At first, I rounded it to 15... but i can't get the answer. So, I changed it to 20

OpenStudy (anonymous):

its why I used 18 instead of 17. Instead of rounding 17.47 directly up to 17, I rounded it up to 17.5, and then rounded up to 18. Then i borrowed numbers already in the problem. 12 loses to and gives it to 18, which makes it 10 + 20 while 6 loses one and gives it to 19, which makes it 20 + 5. all together its 10+20+20+5

OpenStudy (anonymous):

Oh... I see your point!!! Thanks a ton!!!!

OpenStudy (anonymous):

no prob :)